How much energy is stored by the electric field between two square plates, 8.3 cm on a side, separated by a 1.9-mm air gap? The charges on the plates are equal and opposite and of magnitude 13 nC .

I get 1.49*10^-6 J and its somehow wrong. HELP PLEASE!!
Answer #1

Here ,

for the capacitance

C = epsilon * Area/distance

C = 8.854 *10^-12 * 0.083^2/0.0019

C = 3.21 *10^-11 F

charge , Q = 13 nC

energy stored by electric field = 0.50 * Q^2/C

energy stored by electric field = 0.50 * (13 *10^-9)^2/(3.21 *10^-11)

energy stored by electric field = 2.63 *10^-6 J

the energy stored by electric field is 2.63 *10-6 J